# A Study on Malicious Browser Extensions in 2025

### v1

@@ Abstract
Index Terms:
Browser Extensions, Malicious Browser Extensions, Attacks due to Extensions.
@@ 1.2 Malicious Browser Extension Based Attacks
MBEs are third-party add-ons that exploit browser APIs to perform unauthorized actions such as data theft, spying, content manipulation, and session hijacking. Increasingly, these extensions, often disguised as productivity tools, enable attacks including keylogging, data exfiltration, malicious code injection, and session hijacking. Early MBEs (2013–2015) primarily focused on basic data theft—keylogging (e.g., HoverZoom, 2013) and activity tracking (e.g., BBC News Reader, 2014)—eroding user trust. By 2016, attackers employed more advanced techniques such as web request interception (e.g., iCalc, 2016) and token theft via phishing (e.g., Viralands, 2016). From 2017 onward, MBEs adopted complex payloads for financial and operational exploitation, including ad fraud (e.g., Web Developer, 2017) and proxy-based traffic manipulation (e.g., Dubbed Copyfish, 2017). Recently, supply chain attacks (e.g., Cyberhaven, 2024) have emerged, where adversaries compromise trusted developer accounts using social engineering and OAuth token abuse to deploy malicious updates. Table I summarizes notable MBE incidents from 2013 to 2024, illustrating the shift from opportunistic attacks to coordinated, high-value campaigns exploiting browser vulnerabilities.
TABLE I: Malicious Extension Based Attacks (2013-2024) [6],[7],[5],[8],[9],[10],[13],[26]
@@
HoverZoom (2013) Keylogging Browse images on websites by hovering. Collecting online form data and selling users’ keystrokes.
@@
Autocopy (2014) Spying Select text and automatically copy to the clipboard. Sends a lot of user data back to its servers.
@@
Marauder’s Map (2015) Spying Plot your friends’ location data from Facebook on a map. A hacker can know if you’re not home, shops you visit frequently, who you spend most time with.
Viralands (2016) Phishing “Verify your age” to access restricted content. Access to Facebook access token; login credentials stolen.
iCalc (2016) Webpage Manipulation Functional Calculator. Creates a proxy and intercept web requests, taking commands and updates from a domain.
Dubbed Copyfish (2017) Mal-Ads Extract text from images, PDFs, videos. Equipped with ad injection capabilities.
@@
Nano Adblocker/Nano Defender (2020) Spying Adblocker. Collected user data and sent it to remote servers.
The Great Suspender (2021) Malware Suspends unused tabs to save memory. Injected malicious code to steal data.
@@
Cyberhaven (2024) Supply Chain Compromised via phishing targeting developer accounts. Distributed malicious versions, stealing Facebook access tokens and bypassing 2FA.
StealthSpy (2024) Spying Disguised as a productivity enhancer. Secretly records browsing history and keylogs sensitive data.
AdSkimmer Pro (2024) Ad Fraud Claimed to block ads. Injected its own ads and skimmed affiliate commissions from legitimate sites.
QuickAccess Helper (2024) Phishing Promised faster access to commonly visited sites. Redirected users to phishing pages that stole credentials.
@@ 4 Malicious Browser Extensions: Threat Landscape in 2025
Malicious Browser extensions represent a growing threat to user data privacy and security. This section explores the various types of threats posed by these extensions in the year 2025 and discusses their implications. The study in this section has been done by the authors in their laboratory environment over a period of 6 months. The researchers were motivated to build a set of extensions over MV3 for chrome and MV2 for firefox and using the existing APIs exposed by Chrome and Mozilla till December 2024 to cause a security or privacy issue to the user that installs the extensions. The extensions developed were tested by another researcher and verified to be working fine over their browsers before including them into the set of extensions discussed in this paper. Due to the reason that such extensions can be misused and to not provide an easy-to-build environment of such extensions to script kiddies, only a portion of the codes of the extension is displayed here and no completed code reference has been hosted at any platform. The researchers have studied past work and extensions and tested through experiments whether the threats raised via researchers in the past are still there or are patched and whether there are possibilities of new, more capable malicious extensions that can be developed with the new set of APIs which are available to third-party developers. We have discussed some important malicious browser extensions that we experimented and tested during our study. We classified various extensions that we studied based on the types of threat they pose into the below given 5 classes:
@@ 5.1 Cookie Stealing Extension
The extension operates by listening for specific messages from the browser, particularly those requesting cookie information. Upon receiving such a request, it extracts the domain from the active tab’s URL and utilizes the chrome.cookies.getAll (for Chrome) or browser.cookies.getAll (for Firefox) [17] API to retrieve all cookies associated with that domain. The extension then logs these cookies, including their names and values, and exfiltrates them to a remote server using a JavaScript Fetch API request. This method allows attackers to gain access to users’ authentication tokens, leading to session hijacking and unauthorized account access. The permissions required for this operation include cookies, activeTab, and storage as shown in Figure 1.
@@ 5.9 Email Inbox Spying by Browser Extensions
TABLE II: Detailed Discussion of Malicious Browser Extensions Created and Their Bypassing Potential
Name Description API Exploited Required Scripts Permissions Lines of Code Chrome Web Store Mozilla Add-on Store
@@
Cookie Stealing Extension Steals cookies, enabling session hijacking and unauthorized access to user accounts. chrome.cookies.getAll, Fetch API Background Script cookies, activeTab, storage 20-30 Not Possible – Requires cookies permission, detected easily Successfully Published on Add-on Store for Mozilla – Can be obfuscated in a sync manager
@@
Keylogger Extension Captures keystrokes to steal sensitive information such as passwords, credit card numbers, and personal messages. chrome.tabs.query, chrome.tabs.sendMessage Content & Background Script activeTab, scripting, storage 70-90 Not Possible – Keyboard event tracking flagged Successfully Published on Add-on Store for Mozilla – If embedded in a typing tool
@@
Screenshot Capture Extension Secretly captures screenshots of the user’s activity and exfiltrates the data to a remote server for malicious purposes. chrome.tabs.captureVisible Tab, Fetch API Background Script tabs, activeTab, scripting, storage 30-40 Not Possible – Screenshots trigger manual review Successfully Published on Add-on Store for Mozilla – Can justify as a productivity tool
@@
History Tracker Extension Surreptitiously monitors and logs browsing history, enabling user profiling and data exfiltration to external servers. chrome.history.search, chrome.history.getVisits Background Script history, tabs, scripting, storage 40-50 Not Possible – Browsing history requires permissions Successfully Published on Add-on Store for Mozilla – If marketed as ”Browsing Insights”
@@
Auto-Like YouTube Videos Extension Manipulates user engagement on YouTube by automatically liking videos, skewing algorithms, and distorting user activity. MutationObserver, document.querySelector Background Script tabs, activeTab, scripting, storage 40-50 Successfully Published on Chrome Web Store – If embedded in focus tools [35] Successfully Published on Add-on Store for Mozilla – Easily hidden in productivity tools[36]
Manipulation of Web Content Extension Alters webpage content to trick users into phishing schemes, redirecting them to malicious websites, or stealing sensitive information. MutationObserver, chrome.scripting.execute Script Content & Background Script scripting, tabs, storage 50-60 Not Possible – DOM modification triggers review Successfully Published on Add-on Store for Mozilla – If disguised as a UI enhancement
@@
Camera Auto On Extension Activates the user’s camera without consent, enabling remote surveillance and privacy violations, and sending recorded video to a backend server. navigator.mediaDevices. getUserMedia, chrome.scripting.execute Script Background Script camera, tabs, scripting 35-45 Not Possible – Direct camera activation blocked Successfully Published on Add-on Store for Mozilla – Possible with delayed triggers
@@
Injecting Advertisement Malicious Extension Injects unwanted ads that could lead to phishing attacks, malware downloads, or disturb user experience. document.createElement, document.body.appendChild Content & Background Script activeTab 30-40 Possible – Can be hidden in UI tweaks Successfully Published on Add-on Store for Mozilla – If positioned as a customization tool [37]
@@
Email Inbox Spying Extension Monitors and manipulates unread email indicators in webmail, enabling surveillance of private communication and exfiltrates email data. document.querySelector, chrome.scripting.execute Script Content & Background Script scripting, storage, activeTab 20-30 Not Possible – Gmail tracking gets flagged Successfully Published on Add-on Store for Mozilla – Can blend into ”Email Organizer”
@@ 5.10 Analysis and Observations
Our research involved creating and testing multiple malicious browser extensions to evaluate their impact on privacy and security. A key finding is that minimal permissions—such as activeTab, scripting, and storage—are sufficient for executing harmful actions, making these tools easy to develop even for low-skilled attackers. For example, the Cookie Stealing and Keylogger Extensions accessed sensitive data like login credentials and cookies, then exfiltrated this information to remote servers. Similarly, the Screenshot Capture and History Tracker Extensions covertly recorded user activity and captured screenshots, providing attackers with detailed insights into browsing behavior. When evaluating the security measures of Chrome and Firefox extension stores, we observed notable differences in their ability to detect and block malicious extensions. Chrome’s Web Store exhibited stricter security policies, flagging high-risk behaviors such as unauthorized cookie access, keystroke logging, and direct DOM manipulation. Extensions attempting to modify browsing history or inject advertisements were often detected and removed during the review process. On the other hand, Firefox’s Add-ons Store demonstrated greater susceptibility to bypass techniques, particularly when malicious behavior was obfuscated within extensions that appeared to provide legitimate functionality. For instance, extensions that manipulated web content, injected ads, or tracked user activity were more likely to pass Firefox’s review when disguised as productivity tools or interface enhancements. Additionally, we successfully created a To-Do List Extension on Mozilla that contained obfuscated malicious code to inject advertisements, which successfully bypassed Mozilla’s review process and remains active on the Add-ons Store [37]. Similarly, we developed a 25-Minute Timer Extension on Mozilla that included hidden functionality for automatically liking YouTube videos, demonstrating that extensions with benign primary functions can effectively disguise malicious intent and evade detection [36]. On Chrome, we created an extension that displayed a pop-up showing YouTube access time while secretly implementing an auto-liking mechanism for YouTube videos. This extension was successfully published on the Chrome Web Store and remains active, highlighting weaknesses in Chrome’s automated review process for behavioral detection [35].
@@
Table II provides a comparative analysis of these extensions, detailing their capabilities, exploited APIs, required permissions, and the likelihood of bypassing security measures in Chrome and Firefox. Although both browsers employ security mechanisms, the deceptive tactics used by these extensions highlight significant gaps in the review process, particularly in Firefox’s ability to detect obfuscated threats. These findings emphasize the need for more rigorous permission validation, improved static and dynamic analysis tools, and enhanced monitoring mechanisms to prevent unauthorized access to user data and mitigate the risks posed by malicious extensions.
@@ 6 Conclusions
Future research should focus on improving real-time behavioral analysis of browser extensions, refining permission models, and establishing industry-wide vetting standards to mitigate these persistent threats. In particular, developing sophisticated anomaly detection techniques and fostering collaboration between browser vendors, security researchers, and policymakers can enhance the effectiveness of existing defenses. Standardizing permission transparency and incorporating automated rollback mechanisms for post-publication updates will be critical steps in strengthening browser extension security. Ultimately, bridging the gap between theoretical security measures and their real-world implementation is essential. Without proactive efforts from the industry and researchers, malicious extensions will continue to exploit vulnerabilities, posing a persistent threat to the digital ecosystem.

### v2

@@ 1.2 Malicious Browser Extension Based Attacks
MBEs are third-party add-ons that exploit browser APIs to perform unauthorized actions such as data theft, spying, content manipulation, and session hijacking. Increasingly, these extensions, often disguised as productivity tools, enable attacks including keylogging, data exfiltration, malicious code injection, and session hijacking. Early MBEs (2013–2015) primarily focused on basic data theft—keylogging (e.g., HoverZoom, 2013) and activity tracking (e.g., BBC News Reader, 2014)—eroding user trust. By 2016, attackers employed more advanced techniques such as web request interception (e.g., iCalc, 2016) and token theft via phishing (e.g., Viralands, 2016). From 2017 onward, MBEs adopted complex payloads for financial and operational exploitation, including ad fraud (e.g., Web Developer, 2017) and proxy-based traffic manipulation (e.g., Dubbed Copyfish, 2017). Recently, supply chain attacks (e.g., Cyberhaven, 2024) have emerged, where adversaries compromise trusted developer accounts using social engineering and OAuth token abuse to deploy malicious updates. Table 1 summarizes notable MBE incidents from 2013 to 2024, illustrating the shift from opportunistic attacks to coordinated, high-value campaigns exploiting browser vulnerabilities.
Table 1: Malicious Extension Based Attacks (2013-2024)
@@
HoverZoom (2013) [13] Keylogging Browse images on websites by hovering. Collecting online form data and selling users’ keystrokes.
@@
Autocopy (2014) [13] Spying Select text and automatically copy to the clipboard. Sends a lot of user data back to its servers.
@@
Marauder’s Map (2015) [13] Spying Plot your friends’ location data from Facebook on a map. A hacker can know if you’re not home, shops you visit frequently, who you spend most time with.
Viralands (2016) [13] Phishing “Verify your age” to access restricted content. Access to Facebook access token; login credentials stolen.
iCalc (2016) [13] Webpage Manipulation Functional Calculator. Creates a proxy and intercept web requests, taking commands and updates from a domain.
Dubbed Copyfish (2017) [13] Mal-Ads Extract text from images, PDFs, videos. Equipped with ad injection capabilities.
@@
Nano Adblocker/Nano Defender (2020) [6] [9] [13] Spying Adblocker. Collected user data and sent it to remote servers.
The Great Suspender (2021) [7] [13] Malware Suspends unused tabs to save memory. Injected malicious code to steal data.
@@
Cyberhaven (2024) [8] [26] [10] Supply Chain Compromised via phishing targeting developer accounts. Distributed malicious versions, stealing Facebook access tokens and bypassing 2FA.
StealthSpy (2024) [8] Spying Disguised as a productivity enhancer. Secretly records browsing history and keylogs sensitive data.
AdSkimmer Pro (2024) [8] Ad Fraud Claimed to block ads. Injected its own ads and skimmed affiliate commissions from legitimate sites.
QuickAccess Helper (2024) [8] Phishing Promised faster access to commonly visited sites. Redirected users to phishing pages that stole credentials.
@@ 4 Malicious Browser Extensions: Threat Landscape in 2025
Malicious Browser extensions represent a growing threat to user data privacy and security. This section explores the various types of threats posed by these extensions in the year 2025 and discusses their implications. The study in this section has been done by the authors in their laboratory environment over a period of 6 months. All extensions were created and tested between January and May 2025. While the Chrome and Firefox versions used were released in late 2024, the extensions were submitted to their respective stores in February and March 2025, ensuring that the findings reflect the 2025 threat landscape. The researchers were motivated to build a set of extensions over MV3 for chrome and MV2 for firefox and using the existing APIs exposed by Chrome and Mozilla till December 2024 to cause a security or privacy issue to the user that installs the extensions. The extensions developed were tested by another researcher and verified to be working fine over their browsers before including them into the set of extensions discussed in this paper. Due to the reason that such extensions can be misused and to not provide an easy-to-build environment of such extensions to script kiddies, only a portion of the codes of the extension is displayed here and no completed code reference has been hosted at any platform. The researchers have studied past work and extensions and tested through experiments whether the threats raised via researchers in the past are still there or are patched and whether there are possibilities of new, more capable malicious extensions that can be developed with the new set of APIs which are available to third-party developers. We have discussed some important malicious browser extensions that we experimented and tested during our study. We classified various extensions that we studied based on the types of threat they pose into the below given 5 classes:
@@ 5.1 Cookie Stealing Extension
The extension operates by listening for specific messages from the browser, particularly those requesting cookie information. Upon receiving such a request, it extracts the domain from the active tab’s URL and utilizes the chrome.cookies. cgetAll (for Chrome) or browser.cookies.getAll (for Firefox) [17] API to retrieve all cookies associated with that domain. The extension then logs these cookies, including their names and values, and exfiltrates them to a remote server using a JavaScript Fetch API request. This method allows attackers to gain access to users’ authentication tokens, leading to session hijacking and unauthorized account access. The permissions required for this operation include cookies, activeTab, and storage as shown in Figure 1.
@@ 5.9 Email Inbox Spying by Browser Extensions
Table 2: Detailed Discussion of Malicious Browser Extensions Created and Their Bypassing Potential
Name Description API Exploited Chrome Web Store Mozilla Add-on Store
@@
Cookie Stealing Extension Steals cookies, enabling session hijacking and unauthorized access to user accounts. chrome.cookies. getAll, Fetch API Not Possible – Requires cookies permission, detected easily Successfully published on Mozilla – Can be obfuscated in a sync manager
@@
Keylogger Extension Captures keystrokes to steal sensitive information such as passwords and personal messages. chrome.tabs.query, chrome.tabs. sendMessage Not Possible – Keyboard event tracking flagged Successfully published on Mozilla – If embedded in a typing tool
@@
Screenshot Capture Extension Secretly captures screenshots of the user’s activity and exfiltrates the data. chrome.tabs. captureVisibleTab, Fetch API Not Possible – Screenshots trigger manual review Successfully published on Mozilla – Can justify as a productivity tool
@@
History Tracker Extension Monitors and logs browsing history, enabling user profiling. chrome.history. search, chrome.history .getVisits Not Possible – Browsing history requires permissions Successfully published on Mozilla – If marketed as “Browsing Insights”
@@
Auto-Like YouTube Videos Extension Automatically likes videos, skewing engagement metrics. MutationObserver, document.query Selector Successfully published on Chrome – If embedded in focus tools[35] Successfully published on Mozilla – Hidden in productivity tools[36]
Manipulation of Web Content Extension Alters page content to facilitate phishing or redirection. MutationObserver, chrome.scripting. executeScript Not Possible – DOM modifications trigger review Successfully published on Mozilla – Disguised as a UI enhancement
@@
Camera Auto-On Extension Activates the user’s camera without consent for surveillance. navigator. mediaDevices. getUserMedia, chrome.scripting. executeScript Not Possible – Direct camera activation blocked Successfully published on Mozilla – Possible with delayed triggers
@@
Injecting Advertisement Extension Injects unwanted ads that may lead to malware or phishing. document. createElement, document.body. appendChild Possible – Can be hidden in UI tweaks Successfully published on Mozilla – As a customization tool[37]
@@
Email Inbox Spying Extension Monitors unread email indicators and exfiltrates data. document. querySelector, chrome.scripting. executeScript Not Possible – Gmail tracking gets flagged Successfully published on Mozilla – Blended into “Email Organizer”
@@ 5.10 Analysis and Observations
Our research involved creating and testing multiple malicious browser extensions to evaluate their impact on privacy and security. A key finding is that minimal permissions—such as activeTab, scripting, and storage—are sufficient for executing harmful actions, making these tools easy to develop even for low-skilled attackers. For example, the Cookie Stealing and Keylogger Extensions accessed sensitive data like login credentials and cookies, then exfiltrated this information to remote servers. Similarly, the Screenshot Capture and History Tracker Extensions covertly recorded user activity and captured screenshots, providing attackers with detailed insights into browsing behavior. When evaluating the security measures of Chrome and Firefox extension stores, we observed notable differences in their ability to detect and block malicious extensions. Chrome’s Web Store exhibited stricter security policies, flagging high-risk behaviors such as unauthorized cookie access, keystroke logging, and direct DOM manipulation. Extensions attempting to modify browsing history or inject advertisements were often detected and removed during the review process. On the other hand, Firefox’s Add-ons Store demonstrated greater susceptibility to bypass techniques, particularly when malicious behavior was obfuscated within extensions that appeared to provide legitimate functionality. For instance, extensions that manipulated web content, injected ads, or tracked user activity were more likely to pass Firefox’s review when disguised as productivity tools or interface enhancements. Additionally, we successfully created a To-Do List Extension on Mozilla that contained obfuscated malicious code to inject advertisements, which successfully bypassed Mozilla’s review process and remains active on the Add-ons Store [37]. Similarly, we developed a 25-Minute Timer Extension on Mozilla that included hidden functionality for automatically liking YouTube videos, demonstrating that extensions with benign primary functions can effectively disguise malicious intent and evade detection [36]. On Chrome, we created an extension that displayed a pop-up showing YouTube access time while secretly implementing an auto-liking mechanism for YouTube videos. This extension was successfully published on the Chrome Web Store and remains active, highlighting weaknesses in Chrome’s automated review process for behavioral detection [35]. Note: All extensions and code snippets described in this section are original implementations developed by the authors solely for research purposes.
@@
Table 2 provides a comparative analysis of these extensions, detailing their capabilities, exploited APIs, required permissions, and the likelihood of bypassing security measures in Chrome and Firefox. Although both browsers employ security mechanisms, the deceptive tactics used by these extensions highlight significant gaps in the review process, particularly in Firefox’s ability to detect obfuscated threats. These findings emphasize the need for more rigorous permission validation, improved static and dynamic analysis tools, and enhanced monitoring mechanisms to prevent unauthorized access to user data and mitigate the risks posed by malicious extensions. While we have not yet submitted formal disclosures to Mozilla and Chrome regarding our test extensions (e.g., auto-like or ad-injecting behaviors), we recognize the importance of responsible disclosure. We plan to report these results to the appropriate browser security teams prior to final publication, in line with ethical research standards.
@@ 6 Conclusions
Future research should focus on improving real-time behavioral analysis of browser extensions, refining permission models, and establishing industry-wide vetting standards to mitigate these persistent threats. In particular, developing sophisticated anomaly detection techniques and fostering collaboration between browser vendors, security researchers, and policymakers can enhance the effectiveness of existing defenses. Standardizing permission transparency and incorporating automated rollback mechanisms for post-publication updates will be critical steps in strengthening browser extension security. While our experiments focused on Chrome and Firefox on Windows/Linux environments, future work will aim to replicate these findings on macOS and alternative browsers such as Safari and Edge to better generalize the threat landscape across platforms. Ultimately, bridging the gap between theoretical security measures and their real-world implementation is essential. Without proactive efforts from the industry and researchers, malicious extensions will continue to exploit vulnerabilities, posing a persistent threat to the digital ecosystem. Based on our experimental analysis, we propose the following concrete defenses: (1) sandboxed testing environments for pre-publication review, where extensions are executed in isolated sessions to detect hidden behavior; (2) integration of machine learning models that analyze extension behavior over time to catch delayed-onset malicious activities; (3) implementation of permission change alerts to notify users when an extension updates and requests new capabilities; and (4) post-publication telemetry-based monitoring to identify anomalous network or DOM behaviors. These measures, grounded in our findings, can greatly strengthen the browser security ecosystem.